Your recently lost a demand in kansas satet court. After reading the fine print you discovered you have the following two options to pay:

a) you will have to do 31 annual payments of 250000 with the first payment being delivered today.

b) you will have to pay 530000 now. In addition beginning one year from today, you will pay 200000 each year for 30 years.

Using a discount rate of 5.85 percent, witch option should you select? Explain answer.
Answer #1
Rate =5.85% a) PV of costs of first option =(1+r)*PMT*((1-(1+r)^-n)/r) =(1+5.85%)*250000*((1-(1+5.85%)^-31)/5.85%)=3747153.00 b) PV of Costs of 2nd option  PV of costs of 2nd option =First  Payment+PMT*((1-(1+r)^-n)/r) =530000+200000*((1-(1+5.85%)^-30)/5.85%) =3327722.40 Option 2 is better because it provides lower PV of Cost